Janet and Marcus work for a landscaping company. Their project involves making a garden with a 4-foot-wide walkway on each side of a 50-foot square piece of land. The plan for the project is shown.
The walkway will be made of concrete that is 6 inches thick. How many cubic feet of concrete will be needed for the job?

Answer by ewatrrr(24785)


Hi,
Area of a square = side*side
Question states***
Area(w/o walkway) = 50ft*50ft = 2500ft^2
Area( with walkway) = 58ft*58ft = 3364ft^2
Area (walkway) = 3364ft^2 - 2500ft^2 = 864ft^2
volume(walkway 6" thick) = 1/2 ft * 864f^2 = 432ft^3
432ft^3 is the cubic feet of concrete needed for the walkway